OpenStudy (anonymous):

Suppose you won a special math lottery and were given the following options: Option A- You are given $10,000 and recieve $1,000 every year thereafter. Option B- You are given $1,000 and recieve 10% more each year. Which option would you choose and why? What if the lottery went bankrupt in five years, are the plans equal? Would you take a one-time payout of $50,000 instead of option A or B? Why or why not?

OpenStudy (bahrom7893):

I would pick B because I would end up with more than i would've if I had picked A

OpenStudy (bahrom7893):

In five years, hmm 10,000+5*1,000=15,000

OpenStudy (bahrom7893):

Trying to figure out what would be the best way to do this..

OpenStudy (bahrom7893):

Okay cant find a shortcut ill just do it the long way..

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

It's just an annual interest rate formula.

OpenStudy (anonymous):

\[A = P(1+r)^y\]

OpenStudy (bahrom7893):

Polpak what is it? I never learned it, im doing 10 percent for each year

OpenStudy (anonymous):

P is the principal, r is the interest rate (.10) and y is the number of years

OpenStudy (bahrom7893):

oh okay thanks a lot! So P is 1000

OpenStudy (bahrom7893):

So according to polpak: A = 1000*(1+0.10)^5

OpenStudy (bahrom7893):

1610.51

OpenStudy (anonymous):

That is correct

OpenStudy (bahrom7893):

So plans would not be equal, the first option is better, u make more.

OpenStudy (bahrom7893):

Would you take a one-time payout of $50,000 instead of option A or B? Why or why not? Depends on how long the lottery would stay alive.

OpenStudy (anonymous):

It would be interesting to see when they were the cutoff point was.

OpenStudy (anonymous):

\[10 + 5y = (1+.1)^y\]

OpenStudy (anonymous):

Looks like about 60 years. I think I'd go for option A. ;)